OpenStudy (anonymous):

Which set of steps can you use to solve three x minus two equals thirteen in two steps? Subtract 2 from both sides. Then, divide both sides by 3. Add 2 to both sides. Then, multiply both sides by 3. Add 2 to both sides. Then, divide both sides by 3. Subtract 2 from both sides. Then, multiply both sides by 3.

OpenStudy (anonymous):

C

OpenStudy (whpalmer4):

\[3x-2 = 13\] Let's try them out: subtract 2 from both sides: \[3x-2-2 = 13-2\]\[3x-4=11\]That didn't help. Scratch choices A and D \[3x-2=13\] Add 2 to both sides: \[3x-2+2=13+2\]\[3x=15\]Okay, that's promising. Multiply both sides by 3: \[3*3x=3*15\]\[9x=45\]That didn't help, scratch choice B. Divide both sides by 3: \[\frac{3x}{3} = \frac{15}{3}\]\[x=5\]That looks like a winner!